OpenStudy (sophadof):

FAN AND MEDAL! Each key on a piano produces a frequency that is 2^1/12 times higher than the frequency of the key immediately to its left. Moving "n" keys to the right of any key increases the frquency of the starting note by a factor 2^n/12. The key corresponding to Concert A has a frequency of 440 Hz. What is the frequency of note D, which is 5 keys to the right of Concert A?

OpenStudy (roadjester):

Is it \[2^{({n/12)}}\] or \[2^n \over 12\]?

OpenStudy (sophadof):

@roadjester I have no idea..

OpenStudy (sophadof):

i think the first one

OpenStudy (roadjester):

You think? O.o

OpenStudy (sophadof):

yes its the first one @roadjester

OpenStudy (briannaontimeforreal):

i agree

OpenStudy (roadjester):

Okay, so to solve this, what you're going to do is this? \[\large 2^{({n \over 12})}=440\] Next, you're going to take the log base 2, of both sides: \[\large {\log_2}2^{({n \over 12})}={\log_2}440\] Doing this allows you to solve for the value of "n". Once you find n, add 5. Then plug that number back into the original equation.

OpenStudy (roadjester):

And you're done.

OpenStudy (sophadof):

uh.. @roadjester

OpenStudy (roadjester):

I lost you?

OpenStudy (briannaontimeforreal):

im confused

OpenStudy (sophadof):

Yeh... My calculator won't do that @roadjester

OpenStudy (roadjester):

Is your calculator a scientific calculator? Need a scientific to do logarithmic calculations.

OpenStudy (briannaontimeforreal):

im confused @sophadof

OpenStudy (sophadof):

yes it is..i also did it on Mathway.com but it won't allow the \[\log_{2} 2^\frac{ n }{ 12 }\]

OpenStudy (sophadof):

it won't let me do the big 2 after the log 2 @roadjester

OpenStudy (roadjester):

umm....that's because you don't "need to". That's the nice thing about logarithmics. \[\log_\cancel{2} \cancel{2}^{\frac n {12}}=\frac n {12}\]

OpenStudy (roadjester):

oops, missed the log part, \[\cancel{\log_2}\]

OpenStudy (roadjester):

The "big 2" and the log base 2 will cancel out

OpenStudy (sophadof):

ok.. but then how do i figure out n? @roadjester

OpenStudy (roadjester):

multiply by 12 on both sides.... \[12 \times \log_2{440}\]

OpenStudy (sophadof):

105.3? @roadjester

OpenStudy (roadjester):

Yep Now add 5, and that's your "new n"

OpenStudy (sophadof):

so 110.3 @roadjester (:??

OpenStudy (roadjester):

yes, now put it back into ur original equation

OpenStudy (sophadof):

okay! Thank you! @roadjester SO MUUUCH